# Expanding the Chaos: Neural Operator for Stochastic (Partial) Differential Equations

### v1

@@ C.3 Some Examples
In score-based generative models [song2020score], the reverse-time dynamics of the Ornstein-Uhlenbeck process are of central importance. Let $X_{t}\in\mathbb{R}^{d}$ be the state, driven by a $d$-dimensional standard Brownian motion $\overline{W}_{t}$ (hence, $m=d$). The reverse-time SDE is given by:
@@ C.4 A WCE of Path Signatures and a Possible Rough-SDENO
Differential equations driven by signals rougher than Brownian motion with Hurst parameter $\tfrac{1}{4}<H<\tfrac{1}{2}$ appear as variants of common SDEs [cheridito_fractional_2003], as well as in finance and parameter estimation. Such equations are defined by first taking the depth-$N$ (Stratonovich) signature of their drivers
@@
and solving the RDE $dY_{t}=f(Y_{t})\,d\mathbf{X}_{t}$ where $\mathbf{X}$ is lift to a geometric $p$-rough path obtained via the signature. This signature can be viewed as a Taylor expansion capturing higher-order oscillations and cross-dimensional relations, and it uniquely characterises the path up to retracings [lyons_differential_1998]. The truncation level $N$ is set by the driver’s roughness and the vector-field algebra: for a Gaussian driver with Hurst $H$, choose any $p>1/H$ and take $N=\lfloor p\rfloor$. If the vector fields commute or are $k$-step nilpotent, the solution depends only on levels up to $k$, so one may take $N=\min\{k,\lfloor p\rfloor\}$ (e.g., $N=2$ for $H>\tfrac{1}{3}$ and $N=3$ for $\tfrac{1}{4}<H\leq\tfrac{1}{3}$).
For centered Gaussian paths, each grade of $T^{\leq N}(\mathbb{R}^{d})$ admits a finite Wiener–Itô decomposition: the grade-$n$ signature component splits into chaos orders $m\in\{n,n-2,\ldots\}$ [ferrucci2023wienerchaosexpansionsignature]. Let $w_{m}$ be the orthogonal projection onto the $m$-th Wiener chaos. For any multi-index of tensor words $(\gamma_{1},\ldots,\gamma_{n})$ with $n\leq N$,
@@
where $\mathcal{Pair}_{n,m}$ is the set of partitions of $\{1,\dots,n\}$ into (nm)/2(n-m)/2 unordered pairs and $m$ singletons. Each kernel $P_{\Pi;s,t}^{\gamma_{1},\ldots,\gamma_{n}}$ is obtained by contracting, for every pair $\{i,j\}\in\Pi$, the elementary time-ordered kernels with the covariance $R^{\gamma_{i},\gamma_{j}}$ (and its derivatives), while singleton slots remain as symmetrised time kernels; thus $P_{\Pi;s,t}^{\gamma_{1},\ldots,\gamma_{n}}\in L^{2}([s,t]^{m})$ is symmetric in its $m$ variables and $\delta^{(m)}$ denotes the $m$-fold Wiener integral. Ferrucci and Cass’s key contribution is an “integrate–out” step for singletons, via Gaussian integration by parts, that replaces raw time variables with combinations of covariance derivatives (e.g., $\tfrac{1}{2}\,\partial_{v}R(v,v)-\partial_{2}R(s,v)$), yielding $L^{2}$-admissible kernels and convergence for $H>\tfrac{1}{4}$ [ferrucci2023wienerchaosexpansionsignature].
@@ D.1 Wiener–chaos and temporal-basis expansions
Let $X_{t}(x)$ be the stochastic solution process of the SPDE (SPDE). As established, it admits a Wiener–chaos expansion:
@@ D.3 Error decomposition, sensitivity, and universal approximation
for all $f\in C^{s}([0,T])$ (see, e.g., bolin2020numerical and references therein). Applying this pointwise in $(x,\omega)$ and integrating gives (55), with $C_{K}$ absorbing the expected Hölder norms. ∎
@@
For a broad class of SPDEs (in particular, equations with affine coefficients as in neufeld2024solving), it is known that the chaos tail decays factorially in the chaos order. Under the assumptions in neufeld2024solving, one obtains bounds of the form
@@
for suitable constants $C_{M},\Gamma_{T}$ depending on the SPDE coefficients and time horizon but independent of $M$. The proof uses the Stroock–Taylor representation of chaos coefficients in terms of Malliavin derivatives and inductive bounds on these derivatives (see neufeld2024solving); we refer to that work for full details.
@@
The universality of FNO as a neural operator has been established in kovachki2023neural. Let $\mathcal{G}_{M,K}:\mathcal{H}\times\mathbb{R}^{N_{M}}\to\mathcal{H}^{K}$ denote the ideal operator mapping an initial condition $\chi_{0}$ and a finite vector of Wick features $\mathcal{W}(\omega,M)$ to the true temporal coefficients $\{C_{k}(x;\omega)\}_{k=1}^{K}$. Our $\mathcal{F}$-SPDENO model, with an FNO backbone, is a specific parametrisation of this general neural-operator framework. Under the regularity assumptions required by kovachki2023neural, the universal approximation theorem guarantees that for any $\varepsilon>0$ there exists an FNO architecture $\mathcal{G}_{\theta}$ with sufficient capacity such that
@@ Appendix E Detailed Formulation and Analysis of SDENOs
WCE of the SDE solution. By Theorem 1, the strong solution $X_{t}\in\mathbb{R}^{d}$ to the SDE (SDE) admits a Wiener–chaos expansion
@@ F.1 Details in Dynamic $\boldsymbol{\Phi}^{4}_{1}$
The dataset for the one-dimensional dynamic $\Phi^{4}_{1}$ model (also known as the Allen-Cahn or Ginzburg-Landau equation) was generated using a custom numerical solver mirroring the implementation in the torchspde [salvi2022neural] library to ensure consistency with established benchmarks. The process involves two main components: generating the driving Q-Brownian motion and numerically solving the SPDE. The spatial domain was set to the one-dimensional torus $\mathbb{T}^{1}=[0,1]$, discretized with a spatial step of $\Delta x=1/128$. The dynamics were simulated over the time interval $[0,T]$ with a final time of $T=0.05$ and a time step of $\Delta t=10^{-3}$. The driving Q-Brownian motion, $W_{t}$, was synthesized using a Karhunen-Loève-like expansion. Specifically, a standard one-dimensional Brownian motion was generated for each of the $N_{x}+1$ spatial grid points. These independent Brownian paths were then projected onto a sinusoidal basis $\{\sin(j\pi x/L)\}_{j=1}^{N_{x}+1}$ to create a spatially correlated noise field. This construction yields a cumulative space-time noise field $W\in\mathbb{R}^{N\times(N_{t}+1)\times(N_{x}+1)}$, where $N$ is the number of trajectories. The global intensity of the noise was controlled by a parameter $\sigma$, which was set to $\sigma=0.1$. The initial conditions, $\chi_{0}$, were generated to be either deterministic (zero field) or stochastic. In the stochastic case, they were sampled from a random field constructed from a sum of sinusoidal functions with decaying amplitudes, controlled by a parameter $\kappa$. This allows for the study of the solution operator’s dependence on both the noise path $W$ and the initial state $\chi_{0}$. It should be noted that $N=1000$ represents the number of samples used for training; in total, we sampled $1200$ samples, resulting in 200 samples for testing.
@@
The space-time noise $W(t,x)$ was generated to be “white” in time and spatially correlated. This construction follows the spectral approximation of a Q-Brownian motion, where the spatial correlation structure is determined by the eigenfunctions of the covariance operator $Q$. The process begins by generating $N_{x}+1$ independent one-dimensional standard Brownian motions, $(\beta^{j}_{t})_{j=0}^{N_{x}}$, one for each discrete spatial point. The spatially correlated noise field $W(t,x)$ is then constructed by projecting these independent Brownian motions onto a deterministic orthonormal basis, which, in this case, consists of sinusoidal functions $W(t,x)=\sum_{j=0}^{N_{x}}\beta^{j}_{t}\phi_{j}(x)$. Here, the basis functions are given by $\phi_{j}(x)=\sqrt{2/L}\sin(j\pi x/L)$, where $L$ is the length of the spatial domain. This corresponds to the Karhunen-Loève expansion of a Q-Brownian motion whose covariance operator $Q$ is approximated by this sinusoidal basis, a method detailed in [lord2014introduction]. See [salvi2022neural] for more detailed descriptions.
@@
We exactly follow the baselines as well as their architectures included in [salvi2022neural, gong2023deep], including NCDE, NCDE-FNO, NRDE [lu2022comprehensive], FNO [li2020fourier], and NSPDE [salvi2022neural]. We note that we did not include the models presented in [gong2023deep, hu2022neural] in our baselines as they are designed specifically to conquer SPDEs with regularity structures. Furthermore, it may be possible to directly insert the noise trajectory into FNO; however, given that such a setting is not very meaningful, and FNO is commonly used to learn the deterministic coefficients in the Fourier domain, we follow salvi2022neural and claim that $(X_{0},W)\mapsto X$ is not applicable to FNO. It is worth noting that, even though such a setting does not apply to FNO, $\mathcal{F}$-SDENO overcomes this limitation by reordering the chaos and feeding FNO with Q-Brownian trajectory’s Wick features, i.e., $\mathcal{W}(x)$, so that the static, time-independent coefficients can be learned inside FNO.
@@
The $\mathcal{F}$-SPDENO in this experiment is conducted with two major modules, (i) the initial computation of the Wick features, i.e., $\mathcal{W}(x)$ and the selected Haar basis, and (ii) a backbone FNO model [li2020fourier] to learn the time-independent coefficients. The $\mathcal{F}$-SPDENO is trained in a supervised learning framework. The models were trained by minimizing the Mean Squared Error (MSE) loss, which corresponds to the empirical $L^{2}$ error over the training samples. We employed the Adam optimizer for its efficiency and stability in training deep neural networks. The initial learning rate was set to (1e-3), and a batch size of (64) was used. The models were trained for a total of (200) epochs on NVIDIA H200 GPUs. To ensure robust convergence and prevent overfitting, a learning rate scheduler was implemented to reduce the learning rate by a factor of (10) whenever the validation loss plateaued for (15) consecutive epochs. It is worth noting that, in our SPDE settings, we did not include the task of $X_{0}\mapsto X$ as this setting is not applicable to our model since our model computes Wick features based on the chaos trajectories.
@@ F.2 Experimental Details for 2D Navier Stokes Equation
We generate the dataset for the 2D Navier-Stokes equation on a two-dimensional torus $\mathbb{T}^{2}=[0,1]^{2}$, discretized with a spatial resolution of $16\times 16$ and $64\times 64$. The procedure is the same as illustrated in [salvi2022neural, gong2023deep, hu2022neural]. The ground truth solution for the vorticity field $X(t,x)$ is obtained by numerically solving the equation using a pseudo-spectral method. The simulation is run for a total duration of $T=15$ seconds with a time step of $\Delta t=10^{-3}$, resulting in 15,000 time steps per trajectory.
@@ F.3 Experimental Details for Diffusion One-step Sampler
This equation reveals that the dynamics of each propagator $u_{\alpha}(t)$ are composed of two parts: (i) a simple, linear term and a deterministic forcing term (acting only on first-order chaos), both of which are analytically straightforward; and (ii) a complex, non-linear coupling term, $\beta(t)\mathbb{E}[s_{\theta}(t,X_{t})\xi_{\alpha}]$, which encapsulates the entire difficulty of the problem. Crucially, this challenging term is determined by the score function $s_{\theta}(t,X_{t})$, which is approximated by the U-Net in the pre-trained diffusion model. Standard U-Net architectures for diffusion are explicitly conditioned on the time step $t$, typically via a learned time embedding [huang2020unet]. This means the U-Net can serve as a network that outputs the solution of time-reversal SDE implicitly, after feeding it with $t$ and Wick features, i.e., $\xi(\omega)$.
@@
We note that for both CIFAR10 and CelebA-HQ datasets, we sample 1 million $d\overline{W}$ trajectories for our model training, followed by the previous work in [zheng2023fast]. More importantly, given that each trajectory contains 1000 steps, it is not possible to train a model with all steps included. For this case, we conducted the uniform sampling to reduce the time steps to 32 [zheng2023fast].
@@
In Figure 11 we show some additional CIFAR10 learning outcomes (without ground truth from the pretrained model), and in Figure 12 we present unconditional sampling results for Celebhq. In addition, we also show the training loss curve for CIFAR10 in Figure 13.

### v2

@@ C.3 Some Examples
In score-based generative models (Song et al., 2020b), the reverse-time dynamics of the Ornstein-Uhlenbeck process are of central importance. Let $X_{t}\in\mathbb{R}^{d}$ be the state, driven by a $d$-dimensional standard Brownian motion $\overline{W}_{t}$ (hence, $m=d$). The reverse-time SDE is given by:
@@ C.4 A WCE of Path Signatures and a Possible Rough-SDENO
Differential equations driven by signals rougher than Brownian motion with Hurst parameter $\tfrac{1}{4}<H<\tfrac{1}{2}$ appear as variants of common SDEs (Cheridito et al., 2003), as well as in finance and parameter estimation. Such equations are defined by first taking the depth-$N$ (Stratonovich) signature of their drivers
@@
and solving the RDE $dY_{t}=f(Y_{t})\,d\mathbf{X}_{t}$ where $\mathbf{X}$ is lift to a geometric $p$-rough path obtained via the signature. This signature can be viewed as a Taylor expansion capturing higher-order oscillations and cross-dimensional relations, and it uniquely characterises the path up to retracings (Lyons, 1998). The truncation level $N$ is set by the driver’s roughness and the vector-field algebra: for a Gaussian driver with Hurst $H$, choose any $p>1/H$ and take $N=\lfloor p\rfloor$. If the vector fields commute or are $k$-step nilpotent, the solution depends only on levels up to $k$, so one may take $N=\min\{k,\lfloor p\rfloor\}$ (e.g., $N=2$ for $H>\tfrac{1}{3}$ and $N=3$ for $\tfrac{1}{4}<H\leq\tfrac{1}{3}$).
For centered Gaussian paths, each grade of $T^{\leq N}(\mathbb{R}^{d})$ admits a finite Wiener–Itô decomposition: the grade-$n$ signature component splits into chaos orders $m\in\{n,n-2,\ldots\}$ (Ferrucci and Cass, 2023). Let $w_{m}$ be the orthogonal projection onto the $m$-th Wiener chaos. For any multi-index of tensor words $(\gamma_{1},\ldots,\gamma_{n})$ with $n\leq N$,
@@
where $\mathcal{Pair}_{n,m}$ is the set of partitions of $\{1,\dots,n\}$ into (nm)/2(n-m)/2 unordered pairs and $m$ singletons. Each kernel $P_{\Pi;s,t}^{\gamma_{1},\ldots,\gamma_{n}}$ is obtained by contracting, for every pair $\{i,j\}\in\Pi$, the elementary time-ordered kernels with the covariance $R^{\gamma_{i},\gamma_{j}}$ (and its derivatives), while singleton slots remain as symmetrised time kernels; thus $P_{\Pi;s,t}^{\gamma_{1},\ldots,\gamma_{n}}\in L^{2}([s,t]^{m})$ is symmetric in its $m$ variables and $\delta^{(m)}$ denotes the $m$-fold Wiener integral. Ferrucci and Cass’s key contribution is an “integrate–out” step for singletons, via Gaussian integration by parts, that replaces raw time variables with combinations of covariance derivatives (e.g., $\tfrac{1}{2}\,\partial_{v}R(v,v)-\partial_{2}R(s,v)$), yielding $L^{2}$-admissible kernels and convergence for $H>\tfrac{1}{4}$ (Ferrucci and Cass, 2023).
@@ D.1 Wiener–chaos and temporal-basis expansions
Let $X_{t}(x)$ be the stochastic solution process of the SPDE (1). As established, it admits a Wiener–chaos expansion:
@@ D.3 Error decomposition, sensitivity, and universal approximation
for all $f\in C^{s}([0,T])$ (see, e.g., Bolin et al. (2020, Section 3.3) and references therein). Applying this pointwise in $(x,\omega)$ and integrating gives (53), with $C_{K}$ absorbing the expected Hölder norms. ∎
@@
For a broad class of SPDEs (in particular, equations with affine coefficients as in Neufeld and Schmocker (2024)), it is known that the chaos tail decays factorially in the chaos order. Under the assumptions in Neufeld and Schmocker (2024, Sections 4 and 6.6), one obtains bounds of the form
@@
for suitable constants $C_{M},\Gamma_{T}$ depending on the SPDE coefficients and time horizon but independent of $M$. The proof uses the Stroock–Taylor representation of chaos coefficients in terms of Malliavin derivatives and inductive bounds on these derivatives (see Neufeld and Schmocker (2024, Propositions 6.3 and 6.5)); we refer to that work for full details.
@@
The universality of FNO as a neural operator has been established in Kovachki et al. (2023, Section 8.3, Theorem 13). Let $\mathcal{G}_{M,K}:\mathcal{H}\times\mathbb{R}^{N_{M}}\to\mathcal{H}^{K}$ denote the ideal operator mapping an initial condition $\chi_{0}$ and a finite vector of Wick features $\mathcal{W}(\omega,M)$ to the true temporal coefficients $\{C_{k}(x;\omega)\}_{k=1}^{K}$. Our $\mathcal{F}$-SPDENO model, with an FNO backbone, is a specific parametrisation of this general neural-operator framework. Under the regularity assumptions required by Kovachki et al. (2023), the universal approximation theorem guarantees that for any $\varepsilon>0$ there exists an FNO architecture $\mathcal{G}_{\theta}$ with sufficient capacity such that
@@ Appendix E Detailed Formulation and Analysis of SDENOs
WCE of the SDE solution. By Theorem 1, the strong solution $X_{t}\in\mathbb{R}^{d}$ to the SDE (2) admits a Wiener–chaos expansion
@@ F.1 Details in Dynamic $\boldsymbol{\Phi}^{4}_{1}$
The dataset for the one-dimensional dynamic $\Phi^{4}_{1}$ model (also known as the Allen-Cahn or Ginzburg-Landau equation) was generated using a custom numerical solver mirroring the implementation in the torchspde (Salvi et al., 2022) library to ensure consistency with established benchmarks. The process involves two main components: generating the driving Q-Brownian motion and numerically solving the SPDE. The spatial domain was set to the one-dimensional torus $\mathbb{T}^{1}=[0,1]$, discretized with a spatial step of $\Delta x=1/128$. The dynamics were simulated over the time interval $[0,T]$ with a final time of $T=0.05$ and a time step of $\Delta t=10^{-3}$. The driving Q-Brownian motion, $W_{t}$, was synthesized using a Karhunen-Loève-like expansion. Specifically, a standard one-dimensional Brownian motion was generated for each of the $N_{x}+1$ spatial grid points. These independent Brownian paths were then projected onto a sinusoidal basis $\{\sin(j\pi x/L)\}_{j=1}^{N_{x}+1}$ to create a spatially correlated noise field. This construction yields a cumulative space-time noise field $W\in\mathbb{R}^{N\times(N_{t}+1)\times(N_{x}+1)}$, where $N$ is the number of trajectories. The global intensity of the noise was controlled by a parameter $\sigma$, which was set to $\sigma=0.1$. The initial conditions, $\chi_{0}$, were generated to be either deterministic (zero field) or stochastic. In the stochastic case, they were sampled from a random field constructed from a sum of sinusoidal functions with decaying amplitudes, controlled by a parameter $\kappa$. This allows for the study of the solution operator’s dependence on both the noise path $W$ and the initial state $\chi_{0}$. It should be noted that $N=1000$ represents the number of samples used for training; in total, we sampled $1200$ samples, resulting in 200 samples for testing.
@@
The space-time noise $W(t,x)$ was generated to be “white” in time and spatially correlated. This construction follows the spectral approximation of a Q-Brownian motion, where the spatial correlation structure is determined by the eigenfunctions of the covariance operator $Q$. The process begins by generating $N_{x}+1$ independent one-dimensional standard Brownian motions, $(\beta^{j}_{t})_{j=0}^{N_{x}}$, one for each discrete spatial point. The spatially correlated noise field $W(t,x)$ is then constructed by projecting these independent Brownian motions onto a deterministic orthonormal basis, which, in this case, consists of sinusoidal functions $W(t,x)=\sum_{j=0}^{N_{x}}\beta^{j}_{t}\phi_{j}(x)$. Here, the basis functions are given by $\phi_{j}(x)=\sqrt{2/L}\sin(j\pi x/L)$, where $L$ is the length of the spatial domain. This corresponds to the Karhunen-Loève expansion of a Q-Brownian motion whose covariance operator $Q$ is approximated by this sinusoidal basis, a method detailed in (Lord et al., 2014). See (Salvi et al., 2022) for more detailed descriptions.
@@
We exactly follow the baselines as well as their architectures included in (Salvi et al., 2022; Gong et al., 2023), including NCDE, NCDE-FNO, NRDE (Lu et al., 2022b), FNO (Li et al., 2020), and NSPDE (Salvi et al., 2022). We note that we did not include the models presented in (Gong et al., 2023; Hu et al., 2022) in our baselines as they are designed specifically to conquer SPDEs with regularity structures. Furthermore, it may be possible to directly insert the noise trajectory into FNO; however, given that such a setting is not very meaningful, and FNO is commonly used to learn the deterministic coefficients in the Fourier domain, we follow (Salvi et al., 2022) and claim that $(X_{0},W)\mapsto X$ is not applicable to FNO. It is worth noting that, even though such a setting does not apply to FNO, $\mathcal{F}$-SDENO overcomes this limitation by reordering the chaos and feeding FNO with Q-Brownian trajectory’s Wick features, i.e., $\mathcal{W}(x)$, so that the static, time-independent coefficients can be learned inside FNO.
@@
The $\mathcal{F}$-SPDENO in this experiment is conducted with two major modules, (i) the initial computation of the Wick features, i.e., $\mathcal{W}(x)$ and the selected Haar basis, and (ii) a backbone FNO model (Li et al., 2020) to learn the time-independent coefficients. The $\mathcal{F}$-SPDENO is trained in a supervised learning framework. The models were trained by minimizing the Mean Squared Error (MSE) loss, which corresponds to the empirical $L^{2}$ error over the training samples. We employed the Adam optimizer for its efficiency and stability in training deep neural networks. The initial learning rate was set to (1e-3), and a batch size of (64) was used. The models were trained for a total of (200) epochs on NVIDIA H200 GPUs. To ensure robust convergence and prevent overfitting, a learning rate scheduler was implemented to reduce the learning rate by a factor of (10) whenever the validation loss plateaued for (15) consecutive epochs. It is worth noting that, in our SPDE settings, we did not include the task of $X_{0}\mapsto X$ as this setting is not applicable to our model since our model computes Wick features based on the chaos trajectories.
@@ F.2 Experimental Details for 2D Navier Stokes Equation
We generate the dataset for the 2D Navier-Stokes equation on a two-dimensional torus $\mathbb{T}^{2}=[0,1]^{2}$, discretized with a spatial resolution of $16\times 16$ and $64\times 64$. The procedure is the same as illustrated in (Salvi et al., 2022; Gong et al., 2023; Hu et al., 2022). The ground truth solution for the vorticity field $X(t,x)$ is obtained by numerically solving the equation using a pseudo-spectral method. The simulation is run for a total duration of $T=15$ seconds with a time step of $\Delta t=10^{-3}$, resulting in 15,000 time steps per trajectory.
@@ F.3 Experimental Details for Diffusion One-step Sampler
Table 4: Comparison of some fast sampling methods on CIFAR-10 in terms of FID-50K, NFE: number of function evaluations.
Method NFE FID
$\mathcal{U}$-SDENO (ours) 1 4.89
Knowledge distillation (Luhman and Luhman, 2021) 1 9.36
Progressive distillation (Salimans and Ho, 2022) 1 9.12
2 4.51
4 3.00
DDIM (Song et al., 2020a) 10 13.36
20 6.84
50 4.67
SN-DDIM (Bao et al., 2022) 10 12.19
DPM-solver (Lu et al., 2022a) 10 4.70
DEIS (Zhang and Chen, 2022) 10 4.17
@@
This equation reveals that the dynamics of each propagator $u_{\alpha}(t)$ are composed of two parts: (i) a simple, linear term and a deterministic forcing term (acting only on first-order chaos), both of which are analytically straightforward; and (ii) a complex, non-linear coupling term, $\beta(t)\mathbb{E}[s_{\theta}(t,X_{t})\xi_{\alpha}]$, which encapsulates the entire difficulty of the problem. Crucially, this challenging term is determined by the score function $s_{\theta}(t,X_{t})$, which is approximated by the U-Net in the pre-trained diffusion model. Standard U-Net architectures for diffusion are explicitly conditioned on the time step $t$, typically via a learned time embedding (Huang et al., 2020). This means the U-Net can serve as a network that outputs the solution of time-reversal SDE implicitly, after feeding it with $t$ and Wick features, i.e., $\xi(\omega)$.
@@
We note that for both CIFAR10 and CelebA-HQ datasets, we sample 1 million $d\overline{W}$ trajectories for our model training, followed by the previous work in (Zheng et al., 2023). More importantly, given that each trajectory contains 1000 steps, it is not possible to train a model with all steps included. For this case, we conducted the uniform sampling to reduce the time steps to 32 (Zheng et al., 2023).
@@
In Figure 10 we show some additional CIFAR10 learning outcomes (without ground truth from the pretrained model), and in Figure 11 we present unconditional sampling results for Celebhq. In addition, we also show the training loss curve for CIFAR10 in Figure 12. In addition, we provide the FID-50K comparison (in CIFAR10) between our $\mathcal{U}$-SDENO and several classic diffusion fast sampling baselines in Table 4. One can see that $\mathcal{U}$-SDENO shows good performance compared to those distillation methods when the evaluation time is identical, i.e., NFE = 1. Although other methods show better results, they usually require higher NFE numbers, which also verifies the effectiveness of leveraging neural operator-based models as a fast diffusion sampler.